OpenStudy (anonymous):

choose a variablbe to isolate and what it is equal to in the system of equations below. 4x=y+4 4y=3x-1 A.4y=4x-1 B.y=4x+3 C.2y=x-3 D.y=4x-4

OpenStudy (anonymous):

probably easiest to isolate \(y\) in the first equation because that only takes one step \[4x=y+4\iff y=4x-4\]

OpenStudy (anonymous):

thanks again!

OpenStudy (anonymous):

yw again i checked out this site http://www.tylershields.com/

OpenStudy (anonymous):

its pretty cool isnt it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

some pictures look familiar, so i guess he is famous, even if i have never heard the name

OpenStudy (anonymous):

he does really cool shoots with celebs. can you help me with this one?choose a variablbe to isolate and what it is equal to in the system of equations below. 16x=2y+18 4y=3x-1 A.y=8x-9 B.x=-8y+9 C.y=8x+9 D.y=x-9

OpenStudy (anonymous):

now i have to think

OpenStudy (anonymous):

oh it is A \[16x=2y+18\] \[16x-18=2y\] \[8x-9=y\] or \[y=8x-9\]

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

yw